solve for supplementay angle when: 12x+5=9x+7, need the degrees for each angle, and keep getting stuck.
Answer by oberobic(2304)

Supplementary angles total 180.
.
12x + 5 = 9x + 7
subtract 9x from both sides
3x + 5 = 7
subtract 5 from both sides
3x = 2
divide both sides by 3
x = 2/3
.
so the supplementary angle must be 180 - 2/3 = 179 1/3.
.
But that would be a very unusual case.
Did you leave something out?